OpenStudy (anonymous):

For the parent function y equals the cube root of x what effect does the values h = -4 have on the graph?

OpenStudy (anonymous):

Generally, cube root or not, the form would be y-k=\[y-k=\sqrt[3]{x-h}\] Therefore, without k it is just \[y=\sqrt[3]{x-h}\] And then if h=-4 \[y = \sqrt[3]{x--4}\] = \[y = \sqrt[3]{x+4}\]

OpenStudy (anonymous):

I could be wrong about this though because of the fact I've never done cube root shifts. I googled it and found that \[y=\sqrt[3]{x-h} + k\] is the proper format for it so that y is alone. So technically, my prior answer was correct.